OpenStudy (anonymous):

Nick's restaurant bill totaled sixty-four dollars and twenty cents. Show how to estimate fifteen percent of the tip. Show and explain each step.

OpenStudy (anonymous):

ask yourself what is 15% of 64.20? use \[\frac{ is }{ of } \frac{ Percent }{ 100 }\] so you would solve by doing \[\frac{ x }{ 64.20 } \frac{ 15 }{ 100 }\] multiply 15 by 64.20 which gives you 963 divide by 100 you get 9.63 so the tip is $9.63

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

ok here you go You want to leave a 15% tip on a meal that cost $64.20. First, convert the 15% to an actual number that can be used in a calculation. For percents,this is always done by simply dividing the percent (in this case 15%) by 100%.So, the conversational term "15%" becomes 15% / 100% = 0.15 in terms of a real mathematical number. Second, you need to find out what 15% of your $64.20 meal cost is.This is always done by multiplying 0.15 by $64.20, or 0.15 x $64.20=$9.63. So, the amount of tip you are going to leave is $9.63. This makes the total cost of your meal (to write on your charge slip or other payment) $64.20 + $9.63 = $73.83.